Two numbers are respectively 30 percent and 40 percent less than a third number. What percent is the second of the first ?
Answer by Edwin McCravy(20060)

Third number = N
First number = N-(30% of N) = N-(0.30N) = 1.00N - 0.30N = 0.70N
Second number = N-(40% of N) = N-(0.40N) = 1.00N - 0.40N = 0.60N

What percent is the second of the first?
What percent is 0.60N of 0.70N?



Edwin